[Brahms-l] Shift report 20060611 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Jun 11 2006 - 23:59:50 EDT
Shifter leader was dp
Shifters were hy


1. We have beam until 11:00 PM. data taking hours = 5.78 hrs (from run16066-16076, the last run 16077 ended with beam dumped at 11pm, with very low statistics). During our shift MCR called several times to do some polarization task and for that we needed to give some break. no major problem found.  

2. A1-XU in T3 is still have the read-out value at around 5200volts, but all the behaviour of T3 shown in supermon sounds normal. <br>
ted has been not working since the beginning of run06 - it will be very nice if somebody can fix it. 

3. From run16067-16079, the D5 field was set wrongly as 1000 - but they were fixed in the database already. Also, the wrong value on the "run plan of the day" webpage was fixed to be 1050. 

4. Hopefully, another store is coming soon.
_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jun 12 00:00:32 2006

This archive was generated by hypermail 2.1.8 : Mon Jun 12 2006 - 00:00:46 EDT